please help me find the answer to this question.
find the slope between the points (-2,-5) and (3,5)
thank you

Answer by stanbon(75887)

find the slope between the points (-2,-5) and (3,5)
---
slope = (5--5)/(3--2) = 10/5 = 2
=================================
Cheers,
Stan H.